# Uncovering the Viral Nature of Toxicity in Competitive Online Video Games

### v1

@@ Abstract
competitive online video games, peer effects, propagation, toxicity.
@@ I Introduction
Previous research suggests toxicity is so deeply ingrained in competitive online video games that it has become normalized [1]. For instance, a recent study found that 86% of players aged 18 to 45 report having experienced harassment while playing [2]. The incidence of toxic behavior in competitive online video games can be accounted for by their inherently competitive nature, amplified by the anonymity of online interactions [3, 4, 5]. Given these games’ massive player bases, even a low rate of toxic behavior translates into thousands of daily incidents, exposing an even higher number of players to such behavior.
The undesirable effects of toxicity on player performance, psychological well-being, satisfaction, and retention are well-documented [6]. For example, players who are frequently targeted by toxic behavior tend to report more symptoms of depression and problematic gaming habits [7, 8]. Furthermore, both victims and perpetrators tend to report heightened levels of anxiety and anger rumination. Marginalized groups, such as LGBTQ+ individuals, people of color, and women, are disproportionately affected by toxicity and its detrimental effects [9, 10, 11, 12, 13, 14]. In esports, toxic behavior has been found to disrupt team coordination and morale, ultimately hindering players’ in-game performance [15]. Lastly, it has been shown that toxicity contributes to player churn and deters new players from joining, suggesting that video game publishers have a vested interest in mitigating toxicity [16].
@@
Peer effects can be theoretically explained through several mechanisms, one of which is conformity to social norms [27]. As inherently social beings, humans tend to avoid deviating from social norms. Accordingly, if toxicity is perceived as consistent with these norms, as suggested by its frequency and players’ stated perceptions, players may feel compelled or encouraged to engage in such behavior. In this context, the actions of a single individual have the transformative power to redefine the norm, making behavior once considered unacceptable seem acceptable [28]. Therefore, it is plausible to expect toxicity to propagate, with one player’s toxic behavior causing the exposed players to engage in similar behavior. Previous research has already provided correlational evidence in support of this conjecture [29, 30].
@@ IV Discussion and Conclusion
In contrast to many (but not all) studies on toxicity in competitive online video games, ours has the advantage of being observational, avoiding the various biases associated with players’ self-reported behavior and perceptions. This approach comes at the cost of greater analytical complexity, as exemplified by our instrumental variable estimation strategy.
@@
In conclusion, we suggest that future research investigate in priority how the spread of toxic speech to a player’s teammates varies across different contexts. Understanding these differences would be invaluable, as it would allow resources to be allocated to the contexts where the virality of toxicity and, by extension, its adverse effects are the most pronounced.

### v2

@@ Abstract
competitive online video games, peer effects, virality, toxicity.
@@ I Introduction
Previous research suggests toxicity is so deeply ingrained in competitive online video games that it has become normalized, with some players rationalizing it as an inherent part of the game culture [1, 2]. For instance, a recent study found that 86% of players aged 18 to 45 report having experienced harassment while playing [3]. The incidence of toxic behavior in competitive online video games can be accounted for by their inherently competitive nature, amplified by the anonymity of online interactions [4, 5, 6]. Given these games’ massive player bases, even a low rate of toxic behavior translates into thousands of daily incidents, exposing an even higher number of players to such behavior.
The undesirable effects of toxicity on player performance, psychological well-being, satisfaction, and retention are well-documented [7]. For example, players who are frequently targeted by toxic behavior tend to report more symptoms of depression and problematic gaming habits [8, 9]. Furthermore, both victims and perpetrators tend to report heightened levels of anxiety and anger rumination. Marginalized groups, such as LGBTQ+ individuals, people of color, and women, are disproportionately affected by toxicity and its detrimental effects [10, 11, 12, 13, 14, 15, 16]. In esports, toxic behavior has been found to disrupt team coordination and morale, ultimately hindering players’ in-game performance [1]. Lastly, it has been shown that toxicity contributes to player churn and deters new players from joining, suggesting that video game publishers have a vested interest in mitigating toxicity [17].
@@
Peer effects can be theoretically justified through several mechanisms, one of which is conformity to social norms [28]. As inherently social beings, humans tend to avoid deviating from social norms. Accordingly, if toxicity is perceived as consistent with these norms, as suggested by its frequency and players’ stated perceptions, players may feel compelled or encouraged to engage in such behavior. In this context, the actions of a single individual have the transformative power to redefine the norm, making behavior once considered unacceptable seem acceptable [29]. Therefore, it is plausible to expect toxicity to propagate, with one player’s toxic behavior causing the exposed players to engage in similar behavior. Previous research has already provided correlational evidence in support of this conjecture [30, 31]. Relatedly, conformity (or lack thereof) has previously been recognized as a factor in why female and other “out-group” players may be targeted by toxic behavior [13, 7].
@@ IV Discussion and Conclusion
Our work expands on and contributes to the literature on toxicity in competitive online video games by uncovering its viral nature. Rather than relying on survey data, as many studies have done, we take a distinct yet complementary approach, focusing on observational causal inference with gameplay data. This approach mitigates the various biases associated with players’ self-reported behavior and perceptions, though it introduces greater analytical complexity. We formulate an estimation strategy to measure the causal effects of exposure to toxicity. This approach could be applied to further document the effects of toxicity on player performance [1], retention [17], and other key outcomes, thereby strengthening previous research in this area.
@@
In conclusion, we suggest that future research investigate in priority how the spread of toxic speech to a player’s teammates varies across different contexts. Understanding these differences would be invaluable, as it would allow resources to be allocated to the contexts where the virality of toxicity and, by extension, its adverse effects are the most pronounced. Exploring forms of toxicity beyond language and speech also offers a compelling direction for future research.